# sharp bottom spectrum and scalar curvature rigidity

### v1

@@ 4. Sharp Bottom spectrum and scalar curvature rigidity
The proof of sharpness is classical, originating from the methods in [Davaux], where the technique is integrated with classical index theory. However, the net characterization for complete noncompact manifolds and the scalar curvature rigidity characterization for cocompact manifolds in Theorem 1.1 and Theorem 1.3 are innovative. To our best knowledge, they have not previously appeared in the geometric analysis literature under the assumption of scalar curvature lower bound. The key element for the rigidity results is the following unique continuation inequality, which is quite technical. Therefore, the proof is deferred to Section 5. Here, we present only the version involving the Dirac operator, which is necessary for our purposes in this section. For a more general case and its proof, see Section 5.
@@
Suppose that $(X^{n},g)$ is an asymptotically hyperbolic and spin manifold with scalar curvature $\mathrm{Sc}_{g}\geq-n(n-1)$, then
$\lambda_{1}({X}^{n},g)\leq\frac{(n-1)^{2}}{4}.$
@@
Suppose that $(X^{n},g)$ is a complete manifold with non-positive sectional curvature $\sec_{g}\leq 0$ and bounded geometry. If $\mathrm{Sc}_{g}\geq-n(n-1)\kappa$, then
$\lambda_{1}(X,g)\leq\frac{(n-1)^{2}}{4}\kappa.$
@@ 5. Unique continuation Theorem on complete Riemannian manifold
Let $Y$ be a discrete subset of $X$. In this section, we will prove the following unique continuation theorem.
@@
Let $(X^{n},g)$ be a complete Riemannian manifold with bounded geometry. Let $Y$ be a discrete net of $X$ and $N_{a}(Y)$ the $a$-neighborhood of $Y$ for some $a>0$. Let $E$ be a vector bundle over $X$ and $P$ a second order elliptic differential operator acting on $E$ satisfying the Gårding inequality in line (5.1). If $P_{\lambda}$ is the spectral projection of $P$ acting on $L^{2}(E)$ with spectrum $\leq\lambda$ and $V_{\lambda}$ is the range of $P_{\lambda}$, then there exists a constant $C_{\lambda}>0$ such that
@@
Theorem 5.1 is essentially motivated by [MR1362555, MR3041662]. It plays an inessential role in the proof of main theorem regarding the scalar curvature rigidity/scalar curvature distribution.
@@ Appendix A Invertible doubles
In the appendix, we prove an invertibility theorem on doubles, as an application of Theorem 5.1, which is of independent interest.
In [Xieyu14], Xie and Yu proved a higher analogue of the relative index theorem. A key ingredient is the vanishing theorem of the index of an elliptic differential operator on the double of a complete manifold, which is proved in [Xieyu14, Remark 5.5]. In this section, we prove a stronger result that the elliptic operator on a double is indeed invertible, based on Theorem 5.1.

### v2

@@ 4. Sharp Bottom spectrum and scalar curvature rigidity
The proof of the sharpness originates from the methods in [Davaux], where the technique is integrated with the classical index theory. However, the net characterization for complete noncompact manifolds and the scalar curvature rigidity characterization for cocompact manifolds in Theorem 1.2 and Theorem 1.4 are innovative. To our best knowledge, they have not previously appeared in the geometric analysis literature under the assumption of scalar curvature lower bound. The key element for the rigidity results is the following unique continuation inequality, which is quite technical. Therefore, the proof is deferred to Section 5. Here, we present only a version involving the Dirac operator, which is necessary for our purposes in this section. See a more general case and its proof in Section 5.
@@
Suppose that $(X^{n},g)$ is an asymptotically hyperbolic and spin manifold with scalar curvature $\mathrm{Sc}_{g}\geq-\kappa$, then
$\lambda_{1}({X}^{n},g)\leq\frac{n-1}{4n}\kappa.$
@@
Suppose that $(X^{n},g)$ is a complete manifold with non-positive sectional curvature $\sec_{g}\leq 0$ and bounded geometry. If $\mathrm{Sc}_{g}\geq-\kappa$, then
$\lambda_{1}(X,g)\leq\frac{n-1}{4n}\kappa.$
@@ 5. Unique continuation Theorem on complete Riemannian manifold
In this section, we will prove the unique continuation theorem as follows.
@@
Suppose that $(X^{n},g)$ is a complete Riemannian manifold with bounded geometry and $Y$ is a discrete net of $X$ and $N_{a}(Y)$ the $a$-neighborhood of $Y$ for some $a>0$. Let $E$ be a vector bundle over $X$ and $P$ a second order elliptic differential operator acting on $E$ satisfying the Gårding inequality in line (5.1). If $P_{\lambda}$ is the spectral projection of $P$ acting on $L^{2}(E)$ with spectrum $\leq\lambda$ and $V_{\lambda}$ is the range of $P_{\lambda}$, then there exists a constant $C_{\lambda}>0$ such that
@@
Theorem 5.1 is essentially motivated by [MR1362555, MR3041662]. It plays an essential role in the proof of main theorem regarding the scalar curvature rigidity/scalar curvature distribution.
@@ Appendix A Invertible doubles
In the appendix, as an application of Theorem 5.1, we will prove an invertibility theorem on doubles. Recall that Xie-Yu prove a higher analogue of the relative index theorem (see [Xieyu14]). The key ingredient is the vanishing theorem of the index of an elliptic differential operator on the double of a complete manifold (see [Xieyu14]*Remark 5.5. In this section, we prove a stronger result that the elliptic operator on a double is indeed invertible using the Theorem 5.1 in this paper.
@@
Appendix B scalar curvature rigidity and almost flat bundles
Recall that Xie–Yu–Wang develop a quantitative $K$-theory for Lipschitz filtered $C^{*}$-algebras, and then prove the existence of almost flat bundles on spaces with finite asymptotic dimension (see [Wang:2021uj]). Let us recall the definition of asymptotic dimension.
Definition B.1 (Gromov, see [Gromov1993asymptotic]).
The asymptotic dimension of a metric space $X$ is the smallest integer $d$ such that for any $r>0$, there exists a uniformly bounded cover $C_{r}=\{U_{i}\}_{i\in I}$ of $X$ for which the $r$-multiplicity of $C_{r}$ is at most $(d+1)$, that is, no ball of radius $r$ in $X$ intersects more than $(d+1)$ members of $C_{r}$.
One of the main theorems of [Wang:2021uj] is as follows.
Theorem B.2 (Xie–Yu–Wang, see [Wang:2021uj]*Theorem 1.5).
Suppose that $(X,g)$ is a geometrically contractible complete Riemannian manifold with bounded geometry and finite asymptotic dimension. let any $\varepsilon>0$, there exists $R_{\varepsilon}>0$ such that, for any $x\in X$, any $\alpha\in K^{*}(X)=K_{*}(C_{0}(X))$ is represented by an $\varepsilon$-Lipschitz matrix-valued function on $X$ that is supported in the $R_{\varepsilon}$-ball centered at $x$.
In this section, we shall prove a different version of rigidity from Theorem 1.2 under the assumption of finite asymptotic dimension as follows.
Theorem B.3.
Suppose that $(X^{n},g)$ is a complete, geometrically contractible Riemannian manifold with bounded geometry and scalar curvature $\mathrm{Sc}_{g}\geq-\kappa$ for some constant $\kappa\geq 0$. If $(X^{n},g)$ has finite asymptotic dimension, then
$\lambda_{1}(X,g)\leq\frac{n-1}{4n}\kappa.$ (B.1)
$\{p\in X:\mathrm{Sc}_{g}(p)\leq-\kappa+\delta\}$
is a net of $({X},g)$.
Proof.
Without loss of generality, we assume that $X$ is even dimensional, while the odd dimensional case is dealt with similarly by considering $X\times\mathbb{R}$.
The first part is a direct corollary of Theorem 1.2, as spaces with finite asymptotic dimension satisfies the coarse Novikov conjecture [Yu]. Now we assume that $\lambda_{1}(X,g)=\frac{n-1}{4n}\kappa$. It suffices to show that there exists $C_{\delta}>0$ such that for any $p_{0}\in X$, there is some point $p$ in the $C_{\delta}$-neighborhood of $p_{0}$ with $\mathrm{Sc}_{g}(p)\leq-\kappa+\delta$.
Since $X$ has bounded geometry, there exists $C>0$ such that $\mathrm{Sc}_{g}\geq-C$ everywhere on $X$. Set
$C_{\delta}\coloneqq R_{\delta_{1}}+\frac{16}{\delta}\sqrt{\frac{-\kappa+C+\delta}{4}}+1,$
where $\delta_{1}=\sqrt{\frac{\delta}{8n^{2}}}$ and $R_{\delta_{1}}$ is as in Theorem B.2. Assume otherwise that every point $p$ in the $C_{\delta}$-neighborhood of $p_{0}$ has $\mathrm{Sc}_{g}(p)>-\kappa+\delta$.
Since $X$ is geometrically contractible, $X$ is automatically spin, and the Dirac operator $D$ defines a non-zero class $[D]$ in the $K$-homology of $X$, and there exists an element $\alpha\in K^{0}(X)$ such that the pairing between $D$ and $\alpha$ is non-zero, see [willett2020higher, Corollary 9.6.12]. By Theorem B.2, $\alpha$ is represented by $[p]-[q]$, where $p,q$ are $\delta_{1}$-Lipschitz $(N\times N)$-matrix-valued projection functions on $X$, and $(p-q)$ is supported in the $R_{\delta_{1}}$-neighborhood of $p_{0}$. Without loss of generality, we may assume that $q=\begin{pmatrix}1_{k}&0\\ 0&0\end{pmatrix}$, hence $\mathrm{rank}(p)=k$.
Now we consider the relative index of $D$ on $X$ twisted by $[p]-[q]$ as in [GromovLawson]. Let $S$ be the spinor bundle over $X$, equipped with a natural $\mathbb{Z}_{2}$-grading operator $\mathscr{E}$. Denote by $c$ the Clifford action on $S$ and $\nabla$ the spinorial connection on $S$. Consider the rank-$k$ vector bundles $E\coloneqq p\mathbb{R}^{N}$ and $F\coloneqq q\mathbb{R}^{N}$. Set
$V=(S\otimes E)\oplus(S\otimes F)$
to be the twisted spinor bundle. Set
$\nabla^{V}=\begin{pmatrix}(\mathrm{id}_{S}\otimes p)(\nabla\otimes\mathrm{id}_{\mathbb{R}^{N}})(\mathrm{id}_{S}\otimes p)&0\\ 0&(\mathrm{id}_{S}\otimes q)(\nabla\otimes\mathrm{id}_{\mathbb{R}^{N}})(\mathrm{id}_{S}\otimes q)\end{pmatrix}$
to be the twisted connection on $V$. For notation simplicity, we omit $\otimes\mathrm{id}$ and $\mathrm{id}\otimes$ if no confusion arises. The Clifford action on $E$ is given by
$c^{V}(v)=\begin{pmatrix}c(v)&0\\ 0&-c(v)\end{pmatrix}$
for any vector field $v$ over $X$, and the $\mathbb{Z}_{2}$-grading on $V$ is given by
$\mathscr{E}^{V}=\begin{pmatrix}\mathscr{E}&0\\ 0&-\mathscr{E}\end{pmatrix}.$
The twisted Dirac operator on $V$ is given by
$D^{V}=\sum_{i=1}^{n}c^{V}(e_{i})\nabla^{V}_{e_{i}}$
for any local orthonormal basis $\{e_{i}\}$ of $X$. In particular, we have
$D^{V}=\begin{pmatrix}pDp&0\\ 0&qDq\end{pmatrix}.$
Let $\rho\colon X\to\mathbb{R}_{\geq 0}$ be a smooth function such that
  • $\rho$ is $\delta/16$-Lipschitz,
  • $\rho\equiv 0$ inside the $R_{\delta_{1}}$-neighborhood of $p_{0}$, and
  • $\rho(x)\equiv\sqrt{\frac{-\kappa+C+\delta}{4}}$ outside the $C_{\delta}$-neighborhood of $p_{0}$.
Since $p=q=\begin{pmatrix}1_{k}&0\\ 0&0\end{pmatrix}$ outside the $R_{\delta_{1}}$-neighborhood of $p_{0}$, the matrix $\begin{pmatrix}0&1_{k}\\ -1_{k}&0\end{pmatrix}$ defines a unitary operator $U$ on $V$ outside the $R_{\delta/2}$-neighborhood of $p_{0}$. Moreover, we have
$\mathscr{E}^{V}U+U\mathscr{E}^{V}=D^{V}U+UD^{V}=0$
outside the $R_{\delta_{1}}$-neighborhood of $p_{0}$.
Set
$\Psi=\rho\cdot U,$
which is a well-defined smooth endomorphism on $V$ over $X$. Set
$B=D^{V}+\Psi.$
By the relative index theorem in [GromovLawson], $B$ is an odd Fredholm operator acting on $V$, and its Fredholm index is equal to the index pairing of $D$ and $\alpha=[p]-[q]$, which is non-zero. Therefore, there exists a non-zero twisted spinor $s\in L^{2}(X,V)$, such that $Bs=0$.
Note that
$0=\|Bs\|^{2}=$ $\int_{X}|D^{V}s|^{2}+|\Psi s|^{2}+\langle(D^{V}\Psi+\Psi D^{V})s,s\rangle$
$\geq$ $\int_{X}|D^{V}s|^{2}+\rho^{2}|s|^{2}-\frac{\delta}{16}|s|^{2},$
where the last inequality is due to
$D^{V}\Psi+\Psi D^{V}=[D,\rho]U=c^{V}(\nabla\rho)U.$
Direct computation shows that
$(pDp)^{2}=p[D,p]^{2}p+pD^{2}p,\text{ and }(qDq)^{2}=qD^{2}q.$
We note that
$[D,p]=\sum_{i=1}^{n}c^{V}(e_{i})\nabla_{e_{i}}p.$
Hence we have
$\|[D,p]\|\leq n\delta_{1},$
which is independent of the matrix size $N$ of $p$.
By the Lichnerowicz formula, we have
$\|D^{V}s\|^{2}=$ $\int_{X}\big{\langle}\begin{pmatrix}pD^{2}p&0\\ 0&qD^{2}q\end{pmatrix}s,s\big{\rangle}+\big{\langle}\begin{pmatrix}p[D,p]^{2}p&0\\ 0&0\end{pmatrix}s,s\big{\rangle}$
$=$ $\int_{X}\langle\begin{pmatrix}p\nabla^{*}\nabla p&0\\ 0&q\nabla^{*}\nabla q\end{pmatrix}s,s\big{\rangle}+\frac{\mathrm{Sc}_{g}}{4}|s|^{2}+\big{\langle}\begin{pmatrix}p[D,p]^{2}p&0\\ 0&0\end{pmatrix}s,s\big{\rangle}$
$\geq$ $\int_{X}|\nabla s|^{2}+\frac{\mathrm{Sc}_{g}}{4}|s|^{2}-\frac{\delta}{8}|s|^{2}$
$\geq$ $\int_{X}|\nabla^{V}s|^{2}+\frac{\mathrm{Sc}_{g}}{4}|s|^{2}-\frac{\delta}{8}|s|^{2},$
where the last inequality is because $\nabla^{V}=\begin{pmatrix}p\nabla p&0\\ 0&q\nabla q\end{pmatrix}$.
To summarize, we have obtain that
$0\geq\|\nabla^{V}s\|^{2}+\int_{X}\Big{(}\frac{\mathrm{Sc}_{g}}{4}+\rho^{2}-\frac{3\delta}{16}\Big{)}|s|^{2}.$
We claim that
$\frac{\mathrm{Sc}_{g}}{4}+\rho^{2}-\frac{3\delta}{16}\geq\frac{-\kappa+\delta/4}{4}$
by the following.
  • Inside the $C_{\delta}$-neighborhood of $p_{0}$, we have $\mathrm{Sc}_{g}\geq-\kappa+\delta$. Hence
    $\frac{\mathrm{Sc}_{g}}{4}+\rho^{2}-\frac{3\delta}{16}\geq\frac{-\kappa+\delta/4}{4}.$
  • Outside the $C_{\delta}$-neighborhood of $p_{0}$, we have $\mathrm{Sc}_{g}\geq-C$ and $\rho^{2}\equiv(-\kappa+C+\delta)/4$ by construction. Hence
    $\frac{\mathrm{Sc}_{g}}{4}+\rho^{2}-\frac{3\delta}{16}\geq\frac{-C}{4}+\frac{-\kappa+C+\delta}{4}-\frac{3\delta}{16}\geq\frac{-\kappa+\delta/4}{4}.$
Therefore, we see that
$\|\nabla^{V}s\|^{2}\leq\frac{\kappa-\delta/4}{4}\|s\|^{2}.$
Note that the proof of Proposition 3.1 is easily generalized to the twisted Dirac operator $D^{V}$ obtained from the connection $\nabla^{V}$. Therefore, we have
$\|\nabla|s|\|^{2}\leq\frac{n-1}{n}\|\nabla^{V}s\|^{2}\leq\frac{n-1}{4n}(\kappa-\delta/4)\|s\|^{2},$
which contradicts that $\lambda_{1}(X,g)=\frac{n-1}{4n}\kappa$. This finishes the proof. ∎
Remark B.4.
Example 1.3 satisfies the assumptions, hence the conclusions of both Theorem B.3 and Theorem 1.2. Though the assumption of Theorem B.3 is stronger than the one of Theorem 1.2, their rigidity conclusions are independent.